OpenStudy (anonymous):

Find the indicted outputs for f(x)=5 to the second power -2x f(0)= f(-1)= f(2)=

OpenStudy (anonymous):

If it was 0, then it would be 1 because anything raised to the 0 power is 1. -1*-2=2, so it would be 5^2=25 -2*2=-4, so 5 will be in the denominator. You will get 1/(5^4)

OpenStudy (anonymous):

So, f(0)= 1 f(-1)= 25 f(2)=?